# Representative Action Selection for Large Action Space Meta-Bandits

### v3

@@ 4 Generalization and Empirical Validation
First, Sections A-B provide additional related works and the regret bound of Algorithm 1, which were omitted from the main body due to page limits.
@@
Second, Sections C to H contain the proofs of all the theorems and lemmas stated both in the main body and in Section B.
@@
Third, Sections I-J present auxiliary tools used throughout the paper.
@@
Fourth, Sections K-M include additional examples and a numerical study of Algorithm 1. This part features a low-dimensional ($n=3$) action space on the unit sphere to investigate the effect of cluster diameter on expected regret, in line with Theorem 6. It also studies two extensions of the varying dependence structure of actions introduced in the main body:
@@ Appendix A Extra Related Works
Epsilon Nets have two standard definitions. The first, geometric definition [37], requires that radius-$\epsilon$ balls centered at net points cover the set. It relates to the covering number and extends to function classes, as in [31]. The second, measure-theoretic definition [24], requires the net to intersect all subsets of sufficiently large measure. The classic $\epsilon$-net algorithm by [14] remains the simplest and most broadly applicable method. Later works aim to reduce net size [26, 27, 25] and address online settings [6].
Expected supremum of Gaussian process for a given set $\mathcal{S}$ refers to the term $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]$. It is an important topic in high-dimensional probability [37]. The sharpest known bounds are due to [34]. and asymptotic convex geometry [29, Chapter 9]
@@ Appendix B Extra Regret Bounds of Algorithm 1
Theorem 10.
Under the same condition of Theorem 7. Let each cluster contains more than one action and satisfies Assumption 5. For the same constant $C$ in Theorem 6 and another absolute constant $c>0$,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\times c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}.\end{split}$
In fact, the Gaussian process assumption can be relaxed to a sub-Gaussian one, where the random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ satisfies the following increment condition:
@@
Let $\gamma_{2}(\mathcal{S})$ denote Talagrand’s chaining functional [34] for a set $\mathcal{S}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, equipped with the Euclidean norm. It is the sharpest known bound that for a sub-Gaussian process indexed by $\mathcal{S}$, there exists a constant $c>0$ such that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq c\gamma_{2}(\mathcal{S})$. Using this functional, we establish a regret bound for our algorithm:
Theorem 11.
@@
Proofs are in Appendices G and H, respectively.
Appendix C Proof of Lemma 4
@@
where the equality is the probability of missing the cluster $r$ for $K$ times, the first inequality uses the definition of $R_{\epsilon}$. The last inequality uses $\log(1-\epsilon)^{K}=K\log(1-\epsilon)$ and the inequality $\log(1-\epsilon)\leq-\epsilon$ for $\epsilon<1$ [35]. Therefore,
@@
where the first inequality uses union bound, the second inequality uses equation 6. The last inequality uses the fact that there could be at most $\lfloor 1/\epsilon\rfloor$ clusters in $R_{\epsilon}$. ∎
Appendix D Proof of Theorem 6
Under Assumption 5, the regret can be decomposed as follows; note that this Lemma is used only to derive lower bounds.
Lemma 12 (Regret decomposition of reference subsets).
@@ Appendix D Proof of Theorem 6
Under Assumption 5, $\operatorname*{Regret}=Y_{\ell}$ where the cluster $r_{\ell}$ contains the optimal action. Hence,
@@
where the middle equality uses that when $a^{*}(\theta)\in r_{\ell}$, $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}=\max_{a\in r_{\ell}}\mu_{a}$ and further with Assumption 5, $\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\geq\mu_{a_{\ell}}$ when $a^{*}(\theta)\in r_{\ell}$. Then,
@@
Lemma 13 (Expectation integral identity).
@@
where the first equality uses integrated tail formula of expectation (cf. Lemma 1.6.1 of [37]). The last equality set $u_{0}:=\epsilon\sqrt{\log c}$. ∎
Lemma 14 (Borell-TIS inequality; Lemma 2.4.7 of [34]).
@@
It means that the size of the fluctuations of $\mathbb{E}\sup_{a\in\mathcal{S}}X_{a}$ is governed by the size of the individual random variables $X_{a}$.
Statement of Theorem 6:
@@
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}.$
If the partition $\mathcal{R}$ satisfies Assumption 5, then
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}.$
@@
where the first equality follows from the definition of regret in equation 1, and the second equality follows from the assumption that $a^{*}(\theta)\in r_{\ell}$. The inequality holds because $a_{\ell}\in\mathcal{A}$, and hence $\max_{a\in\mathcal{A}}\mu_{a}\geq\mu_{a_{\ell}}$. The final equality follows from the definition of $Y_{\ell}$. On the other hand, according to Lemma 12, under Assumption 5, $\operatorname*{Regret}=Y_{\ell}$, and thus $\operatorname*{Regret}\geq\min_{\ell^{\prime}\leq m}Y_{\ell^{\prime}}$, where the cluster $r_{\ell}$ contains the optimal action.
@@
where the left inequality holds under Assumption 5. Further, by definition $r_{\ell}\subset B(a_{\ell},\epsilon)$, such that $\mathbb{E}Z_{a}^{2}=\mathbb{E}(\mu_{a}-\mu_{a_{\ell}})^{2}=\|a-a_{\ell}\|^{2}_{2}\leq\epsilon^{2}$. Using Lemma 14 on the process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$, we have
@@
Using Lemma 13, we have for some absolute constant $C>0$
@@
where the first inequality uses equation 7. The second inequality uses $\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\leq\mathbb{E}Y_{\ell}+\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses equation 8 and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
@@
where the first inequality uses equation 7 under Assumption 5. The second inequality uses $\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\geq\mathbb{E}Y_{\ell}-\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses equation 8 and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$. ∎
Appendix E Proof of Theorem 7
@@
Lemma 15 (Transformation invariance).
@@
Statement of Theorem 7:
Let $\mathcal{A}$ be output of Algorithm 1. Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Then, for the same constant $C>0$ in Theorem 6,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
@@
If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, define $a_{\ell}\in r_{\ell}\cap\mathcal{A}$. If $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative. The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1. The cluster $r_{\ell}$ is contained in a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$. Define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a random variable
@@
where the first inequality uses $\operatorname*{Regret}\leq Y_{\ell}$ if $r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset$, the equality uses that $Y_{\ell}$ is independent of $r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset$. On the other hand, we assume that $0\in\operatorname*{Conv}(\mathcal{A})$ because even if it doesn’t hold, we can always find a vector $x\in\mathbb{R}^{n}$ such that $0\in\operatorname*{Conv}(\mathcal{A}+x)$, without changing the value of $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ (c.f. Lemma 15). Therefore,
@@
where the first equality uses tower rule. The first inequality uses Equations (11-12), and Lemma 16. The last equality uses tower rule again. The last inequality uses Cauchy-Schwarz inequality, which states that for any two random variables $X$ and $Y$, we have $|\mathbb{E}[XY]|\leq\sqrt{\mathbb{E}[X^{2}]\mathbb{E}[Y^{2}]}$.
Note that the Gaussian process assumption is not used in equation 13; it is only needed to bound the term $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$. By applying equation 9 in the proof of Theorem 6, we bound this term and complete the proof. ∎
Appendix F Proof of Theorem 9
To analyze the worst-case behavior of Theorem 7, we first examine the term $\mathbb{E}_{q}\left[(1-q(r))^{K}\right]$ in the upper bound, summarized in the following Lemma:
Lemma 17.
@@ Appendix F Proof of Theorem 9
When $|\mathcal{R}|\geq K+1$, the maximum is $M=\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$, and it is attained when $q$ is uniform. When $|\mathcal{R}|<K+1$, the maximum is upper bounded by $M\leq\frac{|\mathcal{R}|}{K+1}\left(\frac{K}{K+1}\right)^{K}$, and there exists a feasible distribution $q^{\prime}$ such that $\mathbb{E}_{q^{\prime}}\left[(1-q^{\prime}(r))^{K}\right]\geq\frac{|\mathcal{R}|-1}{K+1}\left(\frac{K}{K+1}\right)^{K}$.
@@
Let $\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be the finite support of the measure $q$, where $m:=|\mathcal{R}|$. Let $q_{i}:=q(r_{i})$ denote the probability mass at each support point. Define the function $f(q_{\ell}):=q_{\ell}\cdot(1-q_{\ell})^{K}$.
@@
Case II $m\geq K+1$: Consider the relaxed maximization problem of equation 14:
@@
To find critical points of equation 15, we solve the system:
@@
Therefore, $q_{\ell}=c$ for all $\ell\leq m$ is a feasible critical point of equation 15. The corresponding objective value is:
@@
$\sum_{\ell=1}^{m}f\left(\frac{1}{m}\right)=\left(1-\frac{1}{m}\right)^{K}=\left(1-\frac{1}{|\mathcal{R}|}\right)^{K},$

### v4

@@ 6 Conclusions and Future Work
First, Sections AC provide additional related work, the regret bound analysis for Algorithm 1, and experimental details omitted from the main text due to space constraints.
@@
Second, Sections D to I contain the proofs of all the theorems and lemmas stated both in the main body and in Section B.
@@
Third, Sections J-K present auxiliary tools used throughout the paper.
@@
Fourth, Sections L-N include additional examples and a numerical study of Algorithm 1. This part features a low-dimensional ($n=3$) action space on the unit sphere to investigate the effect of cluster diameter on expected regret, in line with Theorem 4.2. It also studies two extensions of the varying dependence structure of actions introduced in the main body:
@@ Appendix A Extra Related Works
Multi-Armed Bandits (Lai & Robbins, 1985; Auer et al., 2002a) is defined by a set of actions (arms), each deliver outcomes that are independently drawn from a fixed and unknown distribution. The decision-maker sequentially selects an action, observes its outcome, and aims to maximize cumulative outcomes over time. The policy is to choose the next action based on the sequence of past plays and outcomes. Popular methods include the UCB (Auer et al., 2002a), TS (Agrawal & Goyal, 2012), and EXP3 (Auer et al., 2002b) for adversarial settings.
Optimal Action Identification focuses on identifying the action with the highest expected outcome in a MAB setting using as few samples as possible (Jamieson & Nowak, 2014; Kaufmann et al., 2016). Popular methods in fixed confidence setting include Action Elimination (Even-Dar et al., 2006; Karnin et al., 2013), UCB, and LUCB, all of which achieve sample complexity within a $\log(|\mathcal{A}_{\operatorname*{full}}|)$ factor of the optimum. In fixed budget setting, there is Successive Halving (Karnin et al., 2013), Successive Reject (Audibert & Bubeck, 2010).
Linear Bandits assumes that the expected outcome of each action depends linearly on the inner product between an unknown parameter vector $\theta$ and the feature representation of an action $a\in\mathcal{A}_{\operatorname*{full}}$ (Auer, 2002; Dani et al., 2008; Rusmevichientong & Tsitsiklis, 2010). This line of work assumes that action feature vectors are known a priori, so the cardinality of the action space is not a limiting factor and may even be infinite. The bound scales as $O(n\sqrt{T})$, where $n$ is the dimension of the feature space.
GP Bandits address the case where explicit feature representations are unavailable, but outcomes are assumed to be correlated according to a known kernel function $k(\cdot,\cdot)$ (Williams & Rasmussen, 2006). Srinivas et al. (2009) model the reward function as a sample from a Gaussian process prior and propose the GP-UCB algorithm, which achieves cumulative regret $O\big(\sqrt{T\,\gamma_{T}\log|\mathcal{A}_{\operatorname*{full}}|}\big)$ for finite action spaces, where $\gamma_{T}$ denotes the maximum information gain after $T$ rounds. The widespread adoption of this approach—both for finite action sets (Valko et al., 2013; Li & Scarlett, 2022) and continuous domains (Chowdhury & Gopalan, 2017) highlights the practicality of assuming that action outcomes are correlated.
Epsilon Nets have two standard definitions. The first, geometric definition (Vershynin, 2018), requires that radius-$\epsilon$ balls centered at net points cover the set. It relates to the covering number and extends to function classes, as in Russo & Van Roy (2013). The second, measure-theoretic definition (Matousek, 2013), requires the net to intersect all subsets of sufficiently large measure. The classic $\epsilon$-net algorithm by Haussler & Welzl (1986) remains the simplest and most broadly applicable method. Later works aim to reduce net size (Pach & Tardos, 2011; Rabani & Shpilka, 2009; Mustafa, 2019) and address online settings (Bhore et al., 2024).
Expected supremum of Gaussian process for a given set $\mathcal{S}$ refers to the term $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]$. It is an important topic in high-dimensional probability (Vershynin, 2018) and asymptotic convex geometry (Rothvoss, 2021, Chapter 9). The sharpest known bounds are due to (Talagrand, 2014). The expectation of regret defined in (1) is indeed the difference between the expected suprema of two Gaussian processes, $\{\mu_{a}\}_{a\in\mathcal{A}{\operatorname*{full}}}$ and $\{\mu_{a}\}_{a\in\mathcal{A}}$. This connection motivates our ideas and analysis around the concept of an $\epsilon$-net, which is a central tool for bounding the expected supremum.
@@ Appendix B Extra Regret Bounds of Algorithm 1
Theorem B.1.
Under the same condition of Theorem 4.5. Let each cluster contains more than one action and satisfies Assumption 4.3. For the same constant $C$ in Theorem 4.2 and another absolute constant $c>0$,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq\left(\min_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right]-C\epsilon\sqrt{\log m}\right)\times c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}.\end{split}$
The Gaussian process assumption can be relaxed to a sub-Gaussian one, where the random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ satisfies the following increment condition:
@@
Let $\gamma_{2}(\mathcal{S})$ denote Talagrand’s chaining functional (Talagrand, 2014) for a set $\mathcal{S}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, equipped with the Euclidean norm. It is the sharpest known bound that for a sub-Gaussian process indexed by $\mathcal{S}$, there exists a constant $c>0$ such that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq c\gamma_{2}(\mathcal{S})$.
Using this functional, we establish a regret bound for our algorithm:
Theorem B.2.
@@
Proofs are in Appendices H and I, respectively.
Appendix C Extra Details of Experiments
In general cases, the decision-maker may not have explicit access to the full structure of the action space, especially in high-dimensional settings. Instead, they are given a list of actions and can observe the expected outcomes of these actions through sampling. We therefore treat the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ as a family of random variables indexed by an abstract set. In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{\operatorname*{full}}$: $f(a):=\mu_{a},\forall a\in\mathcal{A}_{\operatorname*{full}}$. We show that our framework applies when the outcome functions lie in a reproducing kernel Hilbert space (RKHS) (Williams & Rasmussen, 2006). Consider a positive semidefinite kernel $k:\mathcal{A}_{\operatorname*{full}}\times\mathcal{A}_{\operatorname*{full}}\to\mathbb{R}$. Consider a positive semidefinite kernel $k:\mathcal{A}_{\operatorname*{full}}\times\mathcal{A}_{\operatorname*{full}}\to\mathbb{R}$. By Mercer’s theorem, for a non-negative measure $\mathbb{P}$ over $\mathcal{A}_{\operatorname*{full}}$, if the kernel satisfies $\int_{\mathcal{A}_{\operatorname*{full}}\times\mathcal{A}_{\operatorname*{full}}}k^{2}(a,a^{\prime})d\mathbb{P}(a)d\mathbb{P}(a^{\prime})<\infty,$ then it admits an eigenfunction expansion: $k(a,a^{\prime})=\sum\nolimits_{i\leq\infty}\lambda_{i}\phi_{i}(a)\phi_{i}(a^{\prime}),$ where $(\phi_{i})_{i\leq\infty}$ are orthonormal eigenfunctions under $\mathbb{P}$, and $(\lambda_{i})_{i\leq\infty}$ are non-negative eigenvalues. Let outcome functionff be of the form $f(\cdot)=\sum_{i=1}^{N}\alpha_{i}k(\cdot,a_{i})$ for some integer $N\geq 1$, and a set of points $\{a_{i}\}_{i=1}^{N}\subset\mathcal{A}_{\operatorname*{full}}$ and a weight vector $\alpha\in\mathbb{R}^{N}$. The function $f$ can be rewritten as
$\mu_{a}=f(a)=\langle\mathbf{f},\Phi(a)\rangle,$
Data Generation:
Assuming the outcome function is sampled from a kernel implies action correlations. To study varying correlation levels, we use the RBF kernel
$k(a,a^{\prime}):=\exp\!\left(-\frac{\|a-a^{\prime}\|^{2}}{2l^{2}}\right),$
where the length-scale $l$ controls dependence among actions. Following (Kanagawa et al., 2018) (Theorem 4.12), we sample functions from the RKHS by forming the kernel matrix $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$ over $\mathcal{A}_{\operatorname*{full}}$ and drawing $f\sim\mathcal{N}(0,\mathbf{K})$. Although the kernel generates the outcomes, it is not revealed to the learner; thus, actions are treated as abstract indices rather than kernel inputs.
Appendix D Proof of Lemma 3.3
@@ Appendix C Extra Details of Experiments
where the equality is the probability of missing the cluster $r$ for $K$ times, the first inequality uses the definition of $R_{\epsilon}$. The last inequality uses $\log(1-\epsilon)^{K}=K\log(1-\epsilon)$ and the inequality $\log(1-\epsilon)\leq-\epsilon$ for $\epsilon<1$ (TOPSØE1, 2007). Therefore,
@@
where the first inequality uses union bound, the second inequality uses (12). The last inequality uses the fact that there could be at most $\lfloor 1/\epsilon\rfloor$ clusters in $R_{\epsilon}$. ∎
Appendix E Proof of Theorem 4.2
Under Assumption 4.3, the regret can be decomposed as follows; note that this Lemma is used only to derive lower bounds.
Lemma E.1 (Regret decomposition of reference subsets).
@@
Under Assumption 4.3, $\operatorname*{Regret}=Y_{\ell}$ where the cluster $r_{\ell}$ contains the optimal action. Hence,
@@
where the middle equality uses that when $a^{*}(\theta)\in r_{\ell}$, $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}=\max_{a\in r_{\ell}}\mu_{a}$ and further with Assumption 4.3, $\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\geq\mu_{a_{\ell}}$ when $a^{*}(\theta)\in r_{\ell}$. Then,
@@
Lemma E.2 (Expectation integral identity).
@@
where the first equality uses integrated tail formula of expectation (cf. Lemma 1.6.1 of (Vershynin, 2018)). The last equality set $u_{0}:=\epsilon\sqrt{\log c}$. ∎
Lemma E.3 (Borell-TIS inequality; Lemma 2.4.7 of (Talagrand, 2014)).
@@
It means that the size of the fluctuations of $\sup_{a\in\mathcal{S}}X_{a}$ is governed by the size of the individual random variables $X_{a}$.
Statement of Theorem 4.2:
@@
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log m}.$
If the partition $\mathcal{R}$ satisfies Assumption 4.3, then
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log m}.$
@@
where the first equality follows from the definition of regret in (1), and the second equality follows from the assumption that $a^{*}(\theta)\in r_{\ell}$. The inequality holds because $a_{\ell}\in\mathcal{A}$, and hence $\max_{a\in\mathcal{A}}\mu_{a}\geq\mu_{a_{\ell}}$. The final equality follows from the definition of $Y_{\ell}$. On the other hand, according to Lemma E.1, under Assumption 4.3, $\operatorname*{Regret}=Y_{\ell}$, and thus $\operatorname*{Regret}\geq\min_{\ell^{\prime}\leq m}Y_{\ell^{\prime}}$, where the cluster $r_{\ell}$ contains the optimal action.
@@
where the left inequality holds under Assumption 4.3. Further, by definition $r_{\ell}\subset B(a_{\ell},\epsilon)$, such that $\mathbb{E}Z_{a}^{2}=\mathbb{E}(\mu_{a}-\mu_{a_{\ell}})^{2}=\|a-a_{\ell}\|^{2}_{2}\leq\epsilon^{2}$. Using Lemma E.3 on the process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$, we have
@@
Using Lemma E.2, we have for some absolute constant $C>0$
@@
where the first inequality uses (13). The second inequality uses $\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\leq\mathbb{E}Y_{\ell}+\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses (14) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
@@
where the first inequality uses (13) under Assumption 4.3. The second inequality uses $\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\geq\mathbb{E}Y_{\ell}-\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses (14) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$. ∎
Appendix F Proof of Theorem 4.5
@@
Lemma F.1 (Transformation invariance).
@@
Statement of Theorem 4.5:
Let $\mathcal{A}$ be output of Algorithm 1. Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Then, for the same constant $C>0$ in Theorem 4.2,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log m}\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
@@
If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, define $a_{\ell}\in r_{\ell}\cap\mathcal{A}$. If $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative. The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 4.1. The cluster $r_{\ell}$ is contained in a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$. Define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a random variable
@@
where the first inequality uses $\operatorname*{Regret}\leq Y_{\ell}$ if $r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset$, the equality uses that $Y_{\ell}$ is independent of $r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset$. On the other hand, we assume that $0\in\operatorname*{Conv}(\mathcal{A})$ because even if it doesn’t hold, we can always find a vector $x\in\mathbb{R}^{n}$ such that $0\in\operatorname*{Conv}(\mathcal{A}+x)$, without changing the value of $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ (c.f. Lemma F.1). Therefore,
@@
where the first equality uses tower rule. The first inequality uses Equations (17-18), and Lemma F.2. The last equality uses tower rule again. The last inequality uses Cauchy-Schwarz inequality, which states that for any two random variables $X$ and $Y$, we have $|\mathbb{E}[XY]|\leq\sqrt{\mathbb{E}[X^{2}]\mathbb{E}[Y^{2}]}$.
Note that the Gaussian process assumption is not used in (19); it is only needed to bound the term $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$. By applying (15) in the proof of Theorem 4.2, we bound this term and complete the proof. ∎
Appendix G Proof of Theorem 4.9
To analyze the worst-case behavior of Theorem 4.5, we first examine the term $\mathbb{E}_{q}\left[(1-q(r))^{K}\right]$ in the upper bound, summarized in the following Lemma:
Lemma G.1.
@@
When $m\geq K+1$, the maximum is $M=\left(1-\frac{1}{m}\right)^{K}$, and it is attained when $q$ is uniform. When $m<K+1$, the maximum is upper bounded by $M\leq\frac{m}{K+1}\left(\frac{K}{K+1}\right)^{K}$, and there exists a feasible distribution $q^{\prime}$ such that $\mathbb{E}_{q^{\prime}}\left[(1-q^{\prime}(r))^{K}\right]\geq\frac{m-1}{K+1}\left(\frac{K}{K+1}\right)^{K}$.
@@
Let $\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be the finite support of the measure $q$, where $m:=m$. Let $q_{i}:=q(r_{i})$ denote the probability mass at each support point. Define the function $f(q_{\ell}):=q_{\ell}\cdot(1-q_{\ell})^{K}$.
@@
Case II $m\geq K+1$: Consider the relaxed maximization problem of (20):
@@
To find critical points of (21), we solve the system:
@@
Therefore, $q_{\ell}=c$ for all $\ell\leq m$ is a feasible critical point of (21). The corresponding objective value is:
@@
$\sum_{\ell=1}^{m}f\left(\frac{1}{m}\right)=\left(1-\frac{1}{m}\right)^{K}=\left(1-\frac{1}{m}\right)^{K},$